OpenStudy (anonymous):

Determine whether the infinite geometric series is convergent or divergent. If it is convergent, find its sum. (If the series is divergent, enter DIVERGENT.)

OpenStudy (anonymous):

\[1+\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }+\frac{ 1 }{ 8 }+...\]

OpenStudy (dbzfan836):

DIVERGENT

OpenStudy (dbzfan836):

It continues, showing 1+1/2+1/4+1/8+1/16+1/32+.... and continues on and on and on...

OpenStudy (dbzfan836):

It keeps adding on, even though they are small, but they still are measurable through equations.

OpenStudy (anonymous):

I tried divergent, but it says that is not correct

OpenStudy (zarkon):

it is a geometric sum that converges

OpenStudy (anonymous):

Thank I already got the answer. 2 \[\frac{ 1 }{ 1-.5 }\]

OpenStudy (zarkon):

2 is the sum